# Physics-informed Value Learner for Offline Goal-Conditioned Reinforcement Learning

### v1

@@ Appendix C Psuedocode and Hyperparameters
The hierarchical policy extraction step follows Park et al. [2024b] and leverages, for both $\pi^{hi}_{\bm{\theta}_{hi}}$ and $\pi^{lo}_{\bm{\theta}_{lo}}$, an advantage-weighted regression-style objective:
@@
where $\beta$ is an inverse temperature hyperparameter and $\tilde{A}^{hi}\left(s_{t},s_{t+k},g\right)$ and $\tilde{A}^{lo}(s_{t},a_{t},s_{t+k})$ are respectively approximated as $V_{\bm{\theta}_{V}}(s_{t+k},g)-V_{\bm{\theta}_{V}}(s_{t},g)$ and $V_{\bm{\theta}_{V}}(s_{t+1},s_{t+k})-V_{\bm{\theta}_{V}}(s_{t},s_{t+k})$. The full pseudocode for Pi-HIQL is provided in Algorithm 1. Furthermore, a function written in JAX on how to compute the gradient $\nabla_{s}V_{\bm{\theta}_{V}}$ in (13) is summarized in Algorithm 2. Finally, Table 3 reports the hyperparameter values most commonly used in our experiments. For more implementation details, refer to our GitHub repository [^2] https://github.com/VittorioGiammarino/PI-HIQL.
Algorithm 1 Physics-informed Hierarchical Implicit Q-Learning (Pi-HIQL)
@@
Table 3: Hyperparameter values for Pi-HIQL.
@@
Appendix D Learning Curves
@@
Figure 10: Learning curves for the experiments in Table Table 2. Plots show the average success percentage per evaluation across seeds as a function of training steps.
Figure 11: Learning curves for the experiments in Table 2. Plots show the average success percentage per evaluation across seeds as a function of training steps.

### v3

@@ Appendix D Psuedocode and Hyperparameters
The hierarchical policy extraction step follows Park et al. [10] and leverages, for both $\pi^{hi}_{\bm{\theta}_{hi}}$ and $\pi^{lo}_{\bm{\theta}_{lo}}$, an advantage-weighted regression-style objective:
@@
where $\beta$ is an inverse temperature hyperparameter and $\tilde{A}^{hi}\left(s_{t},s_{t+k},g\right)$ and $\tilde{A}^{lo}(s_{t},a_{t},s_{t+k})$ are respectively approximated as $V_{\bm{\theta}_{V}}(s_{t+k},g)-V_{\bm{\theta}_{V}}(s_{t},g)$ and $V_{\bm{\theta}_{V}}(s_{t+1},s_{t+k})-V_{\bm{\theta}_{V}}(s_{t},s_{t+k})$. The full pseudocode for Eik-HIQL is provided in Algorithm 1. Furthermore, a function written in JAX on how to compute the gradient $\nabla_{s}V_{\bm{\theta}_{V}}$ in (14) is summarized in Algorithm 2. Finally, Table 3 reports the hyperparameter values most commonly used in our experiments. For more implementation details, refer to our GitHub repository [^2] https://github.com/VittorioGiammarino/Eik-HIQL.
Algorithm 1 Eikonal-regularized Hierarchical Implicit Q-Learning (Eik-HIQL)
@@
Table 3: Hyperparameter values for Eik-HIQL.
@@
Appendix E Learning Curves
@@
Figure 10: Learning curves for the antsoccer experiments in Table 2. Plots show the average success percentage per evaluation across seeds as a function of training steps.
Figure 11: Learning curves for the manipulation experiments in Table 2. Plots show the average success percentage per evaluation across seeds as a function of training steps.
Appendix F Additional Experiments
In the following, we present additional experiments demonstrating that our Eikonal regularizer can be seamlessly integrated with a broad range of temporal-difference (TD)-based GCRL algorithms. In particular, we apply it to Goal-Conditioned variants of IQL [23] and IVL [39, 40], yielding Eik-GCIQL and Eik-GCIVL, respectively. The corresponding results are summarized in Table 4, with learning curves provided in Fig. 12 and 13. These experiments confirm the same conclusions drawn in the main paper from the comparison between Eik-HIQL and HIQL (Table 2), and further support our claim that the Eikonal regularizer can be successfully combined with diverse TD-based algorithms.
Table 4: Summary of the experiments with different TD-based GCRL algorithms. All agents are trained for 100,000 training steps using 10 seeds. We report the mean and standard deviation across seeds for the best evaluation achieved during training. For each seed, evaluations are conducted over $5$ different random goals, as designed in Park et al. [11], with the learned policy tested for $50$ episodes per goal. Results within $95\%$ of the best value are written in bold.
Environment Dataset Type Maze Dimension GCIQL Eik-GCIQL GCIVL Eik-GCIVL
pointmaze navigate medium $60\pm 1$ $59\pm 9$ $63\pm 6$ $90\pm 5$
large $39\pm 1$ $60\pm 9$ $38\pm 5$ $82\pm 39$
giant $0\pm 0$ $2\pm 4$ $0\pm 0$ $86\pm 11$
teleport $29\pm 5$ $25\pm 12$ $38\pm 5$ $49\pm 4$
stitch medium $41\pm 11$ $56\pm 6$ $57\pm 9$ $95\pm 4$
large $5\pm 8 22$$2\pm 3$ $11\pm 8$ $67\pm 9$
giant $\pm 0 0±0$$\pm 0$ $0\pm 0$ $23\pm 10$
teleport $28\pm 5$ $25\pm 3$ $41\pm 5$ $38\pm 3$
antmaze navigate medium $27\pm 4$ $25\pm 6$ $36\pm 5$ $50\pm 5$
large $9\pm 3$ $7\pm 2$ $16\pm 4$ $15\pm 3$
giant $\pm 0 0±0$$\pm 0$ $\pm 0 0±0$$\pm 0$
teleport $24\pm 3$ $23\pm 2$ $32\pm 5$ $30\pm 3$
stitch medium $19\pm 4$ $21\pm 5$ $25\pm 4$ $27\pm 6$
large $6\pm 3$ $3\pm 3$ $12\pm 3$ $7\pm 2$
giant $\pm 0 0±0$$\pm 0$ $\pm 0 0±0$$\pm 0$
teleport $18\pm 5$ $23\pm 3$ $30\pm 3$ $28\pm 3$